The sum of three consecutive integers is 42.
Answer by josmiceli(19441)

let a be the first integer
a + 1 is the next integer
a + 2 is the third
a + (a + 1) + (a +2) = 42
add the integers
3a +3 = 42
combine terms
3a = 39
a = 13
a + 1 = 14
a + 2 = 15
check: 13 + 14 + 15 = 42